# Lepton seesaw model in a modular $A_{4}$ symmetry

### v1

@@ II.2 Neutral fermion mass matrix
$\supset\delta_{N}\overline{N_{L}^{C}}N_{L}+\delta_{\bar{N}}\overline{N_{R}}N_{R}^{C}+m^{\prime}\overline{E_{R}}\ell_{L}+M_{L^{\prime}}\overline{E^{\prime}_{R}}E_{L},$ (35)
@@
and $\delta_{\bar{N}}$ is three by three symmetric matrix with complex values. Note here that $\delta_{\bar{N}}$ does not appear in the neutrino mass matrix as can be seen later. Thus, we will not explicitly write it down.
When we impose the following mass hierarchies $\delta_{N}\ll m^{\prime}\ll M_{L^{\prime}}$, one finds the inverse neutrino mass matrix as follows:
@@
As we can see, $M_{L^{\prime}},\ m^{\prime}$ also contributes to the charged-lepton mass matrix. Furthermore, the mass hierarchies $\delta_{N}\ll m^{\prime},M_{L^{\prime}}$ would be appropriate assumption since $\delta_{N}$ is generated via $v_{\Delta}$ that is restricted by a few GeV to satisfy the constraint on the $\rho$ parameter Zyla et al. (2020). The mass eigenvalues for the active neutrinos $D_{\nu}=\{D_{\nu_{1}},D_{\nu_{2}},D_{\nu_{3}}\}$ are obtained by diagonalizing the mass matrix as $D_{\nu}\equiv V_{\nu}^{T}m_{\nu}V_{\nu}$. Therefore, we obtain $V_{\nu}^{\dagger}m_{\nu}^{\dagger}m_{\nu}V_{\nu}={\rm diag.}(|D_{\nu_{1}}|^{2},|D_{\nu_{2}}|^{2},|D_{\nu_{3}}|^{2})$ where $D_{\nu_{1,2,3}}$ are the three neutrino mass eigenvalues. The Pontecorvo-Maki-Nakagawa-Sakata (PMNS) mixing matrix $U(\equiv U_{PMNS})$ is thus defined by $U=V_{L}^{\dagger}V_{\nu}$ where $V_{L}$ comes from diagonalization of charged-lepton mass matrix. For convenience, we here define dimensionless neutrino mass matrix; $m_{\nu}\equiv\kappa\tilde{m}_{\nu}$ and $\kappa\equiv\frac{v_{\Delta}y_{\Delta}}{\sqrt{2}}$ being a flavor independent mass dimensional parameter. Therefore, $\kappa$ can be rewritten in terms of rescaled neutrino mass eigenvalues $\tilde{D}_{\nu}(\equiv D_{\nu}/\kappa)$ and atmospheric neutrino mass-squared difference $\Delta m_{\rm atm}^{2}$ as follows:
@@
This value should be within the allowed range of the experimental value. Later in numerical analysis, we will adopt NuFit 5.2 Esteban et al. (2020) to experimental ranges of neutrino observables. The effective mass for neutrinoless double beta decay can be written by
$\langle m_{ee}\rangle=\kappa|\tilde{D}_{\nu_{1}}\cos^{2}\theta_{12}\cos^{2}\theta_{13}+\tilde{D}_{\nu_{2}}\sin^{2}\theta_{12}\cos^{2}\theta_{13}e^{i\alpha_{2}}+\tilde{D}_{\nu_{3}}\sin^{2}\theta_{13}e^{i(\alpha_{3}-2\delta_{CP})}|.$ (46)
A predicted value of $\langle m_{ee}\rangle$ is constrained by the current KamLAND-Zen data and could be measured in future Abe et al. (2024). Currently the upper bound is found as $\langle m_{ee}\rangle<(36-156)$ meV at 90 % confidence level where the range of the bound is due to the use of different method in estimating nuclear matrix elements. Furthermore the sum of neutrino masses is constrained by the minimal standard cosmological model $\Lambda$CDM $+\sum D_{\nu_{i}}$ that provides the upper bound $\sum D_{\nu}\leq$ 120 meV Vagnozzi et al. (2017); Aghanim et al. (2020), although it becomes weaker if the data are analyzed in the context of extended cosmological models Olive et al. (2014). Recently, DESI and CMB data combination provides more stringent upper bound on the sum as $\sum D_{\nu}\leq$ 72 meV Adame et al. (2024). These two observable $\langle m_{ee}\rangle$ and $\sum D_{\nu}$ are also discussed as our prediction in the numerical analysis. Direct search for neutrino mass is done by the Karlsruhe Tritium Neutrino (KATRIN) experiment Aker et al. (2022), which is the first sub-eV sensitivity on $m_{\nu_{e}}^{2}=(0.26\pm 0.34)$ eV2 at 90 % CL. Here, $m_{\nu_{e}}^{2}\equiv\sum_{i}|(U_{\rm PMNS})_{ei}|^{2}D^{2}_{\nu_{i}}$. We also take this experiment in account in our numerical analysis below.
III Numerical analysis and phenomenology
In this section, we perform chi square numerical analysis to fit the neutrino data and discuss several features of our model. Furthermore, we concentrate on the region at nearby two representative fixed points $\tau=\{i,\omega\}$ and search for the allowed region within the range of $5\sigma$ with five degrees of freedom. However, we would not find any solutions to satisfy the neutrino oscillation data in case of $\tau=i$. Thus, we focus on discussing features in case of $\tau=\omega$ below.
We randomly select whole our rescaled input parameters within the following ranges $[10^{-3}-10^{3}]$ except for $|\tilde{y}^{\prime}_{\Delta}|$ that is $[10^{-5}-1]$ through our numerical analysis.

### v2

@@ II.2 Neutral fermion mass matrix
$\supset\delta_{N}\overline{N_{L}^{C}}N_{L}+\delta_{\bar{N}}\overline{N_{R}}N_{R}^{C}+m^{\prime}\overline{N_{R}}\nu_{L}+M_{L^{\prime}}\overline{N_{R}}N_{L},$ (40)
@@
and $\delta_{\bar{N}}$ is three by three symmetric matrix with complex values. Note here that $\delta_{\bar{N}}$ does not appear in the neutrino mass matrix as can be seen later. Thus, we will not explicitly write it down. The resultant neutral fermion mass matrix in basis of $\nu_{L},N_{L},N^{C}_{R}]^{T}$ is found as follows:
$\left(\begin{array}[]{ccc}0&0&m^{\prime T}\\ 0&\delta_{N}&M_{L^{\prime}}^{T}\\ m^{\prime}&M_{L^{\prime}}&\delta_{\bar{N}}\end{array}\right).$ (51)
When we impose the following mass hierarchies $\delta_{N}\ll m^{\prime}\ll M_{L^{\prime}}$, one finds the neutrino mass matrix via inverse seesaw as follows:
@@
As we can see, $M_{L^{\prime}},\ m^{\prime}$ also contributes to the charged-lepton mass matrix. Here we comment on our assumption regarding hierarchy of mass parameters; $\delta_{N}\ll m^{\prime}\ll M_{L^{\prime}}$. This can be appropriate assumption as follows. Firstly $\delta_{N}$ is generated via $v_{\Delta}$ that is restricted to be smaller than a few GeV to satisfy the constraint on the $\rho$ parameter Zyla and others (2020), and it can be naturally small due to type-II seesaw mechanism. The mass scale of $m^{\prime}$ is determined by VEV $v^{\prime}$ which would not be far from the electroweak scale where we would choose TeV scale. Finally, scale of $M_{L^{\prime}}$ is free and can be much larger than the electroweak scale. Note also that large $M_{L^{\prime}}$ scale ($\mathcal{O}(100)$ TeV) is motivated to avoid constraints from LFV decays of charged leptons as discussed below.
The mass eigenvalues for the active neutrinos $D_{\nu}=\{D_{\nu_{1}},D_{\nu_{2}},D_{\nu_{3}}\}$ are obtained by diagonalizing the mass matrix as $D_{\nu}\equiv V_{\nu}^{T}m_{\nu}V_{\nu}$. Therefore, we obtain $V_{\nu}^{\dagger}m_{\nu}^{\dagger}m_{\nu}V_{\nu}={\rm diag.}(|D_{\nu_{1}}|^{2},|D_{\nu_{2}}|^{2},|D_{\nu_{3}}|^{2})$ where $D_{\nu_{1,2,3}}$ are the three neutrino mass eigenvalues. We adopt the standard parametrization of the Pontecorvo-Maki-Nakagawa-Sakata (PMNS) mixing matrix $U$ defined by $V_{e_{L}}^{\dagger}V_{\nu}$ where the Majorana phase is defined by $1,e^{i\alpha_{21}/2},e^{i\alpha_{31}/2}]$ Okada and Tanimoto (2021c). Then, the Majorana phases are found as
$\cos\left(\frac{\alpha_{21}}{2}\right)=\frac{\text{Re}[U^{*}_{e1}U_{e2}]}{c_{12}s_{12}c_{13}^{2}},\ \cos\left(\frac{\alpha_{31}}{2}-\delta_{CP}\right)=\frac{\text{Re}[U^{*}_{e1}U_{e3}]}{c_{12}s_{13}c_{13}},$ (53)
$\sin\left(\frac{\alpha_{21}}{2}\right)=\frac{\text{Im}[U^{*}_{e1}U_{e2}]}{c_{12}s_{12}c_{13}^{2}},\ \sin\left(\frac{\alpha_{31}}{2}-\delta_{CP}\right)=\frac{\text{Im}[U^{*}_{e1}U_{e3}]}{c_{12}s_{13}c_{13}},$ (54)
where $\alpha_{21}/2,\ \frac{\alpha_{31}}{2}-\delta_{CP}$ are subtracted from $\pi$, when $\cos(\alpha_{21}/2)$ and $\cos\left(\frac{\alpha_{31}}{2}-\delta_{CP}\right)$ are negative. Similar to Majorana phases, $\delta_{CP}$ is given by
$\cos\delta_{CP}$ $=\frac{-1}{2c_{12}s_{12}c_{23}s_{23}s_{13}}(|U_{31}|^{2}-s_{12}^{2}s_{23}^{2}-c_{12}^{2}c_{23}^{2}s_{13}^{2}),$ (55)
$\sin\delta_{CP}$ $=\frac{1}{s_{23}c_{23}s_{12}c_{12}s_{13}c_{13}^{2}}{\rm Im}[U_{11}U_{22}U_{12}^{*}U_{21}^{*}].\$(56)
For convenience, we define dimensionless neutrino mass matrix $m_{\nu}\equiv\kappa\tilde{m}_{\nu}$ where $\kappa\equiv\frac{v_{\Delta}y_{\Delta}}{\sqrt{2}}$ is a flavor independent mass dimensional parameter. Therefore, $\kappa$ can be rewritten in terms of rescaled neutrino mass eigenvalues $\tilde{D}_{\nu}(\equiv D_{\nu}/\kappa)$ and atmospheric neutrino mass-squared difference $\Delta m_{\rm atm}^{2}$ as follows:
@@
This value should be within the allowed range of the experimental value. In numerical analysis, we will adopt NuFit 5.2 Esteban et al. (2020) to experimental ranges of neutrino observables. The effective mass for neutrinoless double beta decay can be written by
$\langle m_{ee}\rangle=\kappa|\tilde{D}_{\nu_{1}}\cos^{2}\theta_{12}\cos^{2}\theta_{13}+\tilde{D}_{\nu_{2}}\sin^{2}\theta_{12}\cos^{2}\theta_{13}e^{i\alpha_{21}}+\tilde{D}_{\nu_{3}}\sin^{2}\theta_{13}e^{i(\alpha_{31}-2\delta_{CP})}|.$ (59)
[…]
$\kappa^{2}\left[(\tilde{D}_{\nu_{1}}c_{13}c_{12})^{2}+(\tilde{D}_{\nu_{2}}c_{13}s_{12})^{2}+(\tilde{D}_{\nu_{3}}s_{13})^{2}\right].$ (60)
We also take this experiment in account in our numerical analysis below.
II.3 Charged Lepton Flavor Violation
Here we briefly discuss charged lepton flavor violation (CLFV) in the model. The CLFVs can be induced via Yukawa interactions which are the first two terms in the RHS of Eq. (2). Dominant contributions come from one-loop diagram generating dipole operator ;
$\mathcal{L}_{\rm eff}=\frac{C_{L}^{ji}}{M^{2}}\frac{v}{\sqrt{2}}\overline{\ell_{j}}\sigma_{\mu\nu}P_{L}\ell_{i}F^{\mu\nu}+\frac{C_{R}^{ji}}{M^{2}}\frac{v}{\sqrt{2}}\overline{\ell_{j}}\sigma_{\mu\nu}P_{R}\ell_{i}F^{\mu\nu},$ (61)
where $F^{\mu\nu}$ is the electromagnetic field strength tensor and $M$ indicate the heavy particle mass inside loop diagrams ($M=M_{L^{\prime}}$ in our case). The $C_{L,R}^{ji}$ is a Wilson coefficient obtained by calculating one-loop diagram. The order of Wilson coefficient is roughly estimated as $C_{L,R}\sim\frac{y^{2}}{(4\pi)^{2}}\frac{m_{\ell_{i}}}{v}I_{\rm loop}$ where $I_{\rm loop}$ is a $\mathcal{O}(1)$ factor from loop integration and $y$ represents a Yukawa coupling inside a loop diagram. We then consider $\mu\to e\gamma$ process that would provide the strongest CLFV constraints. From the effective interaction, we can estimate
$Br(\mu\to e\gamma)\sim$ $\frac{12\sqrt{2}\pi}{G_{F}^{2}M^{4}}\left[\frac{y^{2}}{(4\pi)^{2}}\frac{m_{\mu}}{v}I_{\rm loop}\right]^{2}$
$\sim$ $10^{-5}y^{4}\left(\frac{\rm TeV}{M}\right)^{4},$ (62)
where $G_{F}$ is the Fermi constant, and we assumed $I_{\rm loop}\sim 1$. Thus we can avoid the current constraint , $Br(\mu\to e\gamma)<1.5\times 10^{-13}$, by choosing $M\gtrsim 100$ TeV even if Yukawa couplings are $\mathcal{O}(1)$. In our numerical analysis, therefore, we assume $M_{L^{\prime}}\gtrsim 100$ TeV and do not discuss CLFVs in detail.
III Mass matrices around the fixed points
Here, we discuss modular forms at nearby fixed points, $\tau=\{i,\ \omega\}$, and show concrete structure of mass and mixing matrices in terms of each fixed point and its deviation.
III.1 $\tau\simeq i$
In case of $\tau\simeq i$, we can expand modular Yukawa couplings in terms of $\epsilon\equiv\tau-i$ Abe et al. (2025). Here, we expand $Y_{3}^{(2)}$ up to $\epsilon$ as follows:
$y_{1}$ $\approx Y_{0}+i\left(\frac{c_{20}}{\sqrt{2}}+\frac{c^{\prime}_{20}}{\sqrt{6}}-\frac{c_{21}}{2\sqrt{3}}\right)\epsilon+{\cal O}(\epsilon^{2}),$ (63)
$y_{2}$ $\approx(1-\sqrt{3})Y_{0}+i\left(\frac{c_{20}}{\sqrt{2}}+\frac{c^{\prime}_{20}}{\sqrt{6}}-\frac{c_{21}}{2\sqrt{3}}\right)\epsilon+{\cal O}(\epsilon^{2}),$ (64)
$y_{3}$ $\approx(-2+\sqrt{3})Y_{0}-i\left(\sqrt{\frac{2}{3}}c^{\prime}_{20}+\frac{c_{21}}{2\sqrt{3}}\right)\epsilon+{\cal O}(\epsilon^{2}),$ (65)
$c_{20}\simeq 1.252$, $c^{\prime}_{20}\approx 0.336$, and $c_{21}\approx 3.049$. Since $m^{\prime}$ as well as $m_{e}$ is represented via $Y_{3}^{(2)}$, the form of charged-lepton mass matrix proportional to
$\left(\begin{array}[]{ccc}y_{1}&y_{3}&y_{2}\\ y_{2}&y_{1}&y_{3}\\ y_{3}&y_{2}&y_{1}\end{array}\right)\simeq Y_{0}\left(\begin{array}[]{ccc}1&-2+\sqrt{3}&1-\sqrt{3}\\ 1-\sqrt{3}&1&-2+\sqrt{3}\\ -2+\sqrt{3}&1-\sqrt{3}&1\end{array}\right)+\frac{i}{2\sqrt{3}}\left(\begin{array}[]{ccc}a_{0}&c_{0}&b_{0}\\ b_{0}&a_{0}&c_{0}\\ c_{0}&b_{0}&a_{0}\end{array}\right)\epsilon+{\cal O}(\epsilon^{2}),$ (75)
where $a_{0}=\sqrt{6}c_{20}+\sqrt{2}c^{\prime}_{20}-c_{21}$, $b_{0}=-\sqrt{6}c_{20}+\sqrt{2}c^{\prime}_{20}-c_{21}$, and $c_{0}=-2\sqrt{2}c^{\prime}_{20}-c_{21}$. Thus, the $V_{e_{L}}$ to diagonalize the dominant matrix, which is denoted by $V_{e_{L}}^{0}$, is found as
$V_{e_{L}}^{0}=\left(\begin{array}[]{ccc}\frac{1}{\sqrt{3}}&\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{6}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{6}}\\ \frac{1}{\sqrt{3}}&0&-\sqrt{\frac{2}{3}}\end{array}\right)\left(\begin{array}[]{ccc}1&0&0\\ 0&\cos\theta&-\sin\theta\\ 0&\sin\theta&\cos\theta\end{array}\right).$ (82)
$\sin\theta$ is determined by solving the following condition:
$(-2+\sqrt{3})(2a_{e}^{2}-b_{e}^{2}-c_{e}^{2})\cos 2\theta+(-3+2\sqrt{3})(b_{e}^{2}-c_{e}^{2})\sin 2\theta=0.$ (83)
At the zeroth order of charged-lepton mass eigenvalues; $m_{e}^{0},m^{0}_{\mu},m^{0}_{\tau}$, we have $m_{e}^{0}=0$ where we abbreviate the other two mass eigenvalues due to too complicated forms. It suggests that the electron mass eigenvalue has to rely on the correction terms $\epsilon$. Hereafter, we redefine $\epsilon_{a}=\frac{a_{e}}{M_{1}}$, $\epsilon_{b}=\frac{b_{e}}{M_{2}}$, and $\epsilon_{c}=\frac{c_{e}}{M_{3}}$.
Then, the neutrino mass matrix at the leading order is given by
$m_{\nu}^{0}\simeq Y_{0}^{2}\frac{v_{\Delta}y_{\Delta}}{\sqrt{2}}\left[\tilde{y}^{\prime}_{\Delta}\epsilon_{b}\epsilon_{c}\left(\begin{array}[]{ccc}-10+6\sqrt{3}&2-\sqrt{3}&8-5\sqrt{3}\\ 2-\sqrt{3}&2-2\sqrt{3}&-4+3\sqrt{3}\\ 8-5\sqrt{3}&-4+3\sqrt{3}&2(-2+\sqrt{3})\end{array}\right)+\left(\begin{array}[]{ccc}1&-2+\sqrt{3}&1-\sqrt{3}\\ -2+\sqrt{3}&7-4\sqrt{3}&-5+3\sqrt{3}\\ 1-\sqrt{3}&-5+3\sqrt{3}&4-2\sqrt{3}\end{array}\right)\epsilon_{a}^{2}\right].$ (90)
The mixing matrix is given by
$U_{\nu}^{0}=\left(\begin{array}[]{ccc}\frac{1}{\sqrt{3}}&\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{6}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{6}}\\ \frac{1}{\sqrt{3}}&0&-\sqrt{\frac{2}{3}}\end{array}\right)\left(\begin{array}[]{ccc}1&0&0\\ 0&\frac{1}{\sqrt{2}}&-\frac{1}{\sqrt{2}}\\ 0&\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{2}}\end{array}\right).$ (97)
Then, the mass eigenvalues at leading order are found as
$D_{\nu_{1}}^{0}=0,\ D_{\nu_{2}}^{0}=9\sqrt{7-4\sqrt{3}}Y_{0}^{2}\tilde{y}^{\prime}_{\Delta}\epsilon_{b}\epsilon_{c},,\ D_{\nu_{3}}^{0}=3\sqrt{7-4\sqrt{3}}Y_{0}^{2}(2\epsilon^{2}+\tilde{y}^{\prime}_{\Delta}\epsilon_{b}\epsilon_{c}).$ (98)
The PMNS at the leading order; $U^{0}$, is given by
$U^{0}=\left(\begin{array}[]{ccc}1&0&0\\ 0&\frac{\cos\theta+\sin\theta}{\sqrt{2}}&\frac{-\cos\theta+\sin\theta}{\sqrt{2}}\\ 0&\frac{\cos\theta-\sin\theta}{\sqrt{2}}&\frac{\cos\theta+\sin\theta}{\sqrt{2}}\end{array}\right).$ (102)
As can be seen from the above form, 1-2 mixing and 1-3 mixing angles have to be generated via correction terms. However 1-2 mixing is not so small via experiments and we could suppose that $\tau=i$ would be disfavored. In fact we have checked that there is no solutions at nearby $\tau=i$ unless $\Delta\chi^{2}$ is greater than $10^{3}$. Thus, we concentrate on the region at nearby $\tau=\omega$.
III.2 $\tau\simeq\omega$
In case of $\tau\simeq\omega$, we can expand modular Yukawa couplings in terms of $\epsilon\equiv\tau-\omega$ Abe et al. (2025). Here, we expand $Y_{3}^{(2)}$ up to $\epsilon$ follows:
$y_{1}$ $\approx\frac{2}{3}\{b_{0}-\frac{i}{\sqrt{3}}(b_{1}-2b_{0})\epsilon+{\cal O}(\epsilon^{2})\},$ (103)
$y_{2}$ $\approx\frac{\omega}{3}\{2b_{0}+\frac{i}{\sqrt{3}}(b_{1}+4b_{0})\epsilon+{\cal O}(\epsilon^{2})\},$ (104)
$y_{3}$ $\approx-\frac{\omega^{2}}{3}\{b_{0}+\frac{2i}{\sqrt{3}}(b_{1}+b_{0})\epsilon+{\cal O}(\epsilon^{2})\},$ (105)
$b_{0}\simeq 1.423$ and $b_{1}\approx 3.673$. Since $m^{\prime}$ as well as $m_{e}$ is represented via $Y_{3}^{(2)}$, the form of charged-lepton mass matrix proportional to
$\left(\begin{array}[]{ccc}y_{1}&y_{3}&y_{2}\\ y_{2}&y_{1}&y_{3}\\ y_{3}&y_{2}&y_{1}\end{array}\right)\simeq\frac{b_{0}}{3}\left(\begin{array}[]{ccc}2&-\omega^{2}&2\omega\\ 2\omega&2&-\omega^{2}\\ -\omega^{2}&2\omega&2\end{array}\right)+\frac{i}{3\sqrt{3}}\left(\begin{array}[]{ccc}2(2b_{0}-b_{1})&-2(b_{0}+b_{1})\omega^{2}&(4b_{0}+b_{1})\omega\\ (4b_{0}+b_{1})\omega&2(2b_{0}-b_{1})&-2(b_{0}+b_{1})\omega^{2}\\ -2(b_{0}+b_{1})\omega^{2}&(4b_{0}+b_{1})\omega&2(2b_{0}-b_{1})\end{array}\right)\epsilon+{\cal O}(\epsilon^{2}).$ (115)
Thus, the $V_{e_{L}}$ to diagonalize the dominant matrix, which is denoted by $V_{e_{L}}^{0}$, is found as
$V_{e_{L}}^{0}=\frac{1}{3}\left(\begin{array}[]{ccc}2\omega&-2\omega&-\omega\\ -\omega^{2}&-2\omega^{2}&2\omega^{2}\\ 2&1&2\end{array}\right).$ (119)
And the zeroth order of charged-lepton mass eigenvalues; $m_{e}^{0},m^{0}_{\mu},m^{0}_{\tau}$, are given by
$m_{e}^{0}=\sqrt{\frac{vv^{\prime}}{2}}\frac{a_{e}f_{1}}{M_{1}}b_{0},\ m_{\mu}^{0}=\sqrt{\frac{vv^{\prime}}{2}}\frac{b_{e}f_{2}}{M_{2}}b_{0},\ m_{\tau}^{0}=\sqrt{\frac{vv^{\prime}}{2}}\frac{c_{e}f_{2}}{M_{3}}b_{0}.$ (120)
It implies that we have the following hierarchy from the experimental masses for the charged-leptons
$\frac{a_{e}}{M_{1}}\ll\frac{b_{e}}{M_{2}}<\frac{c_{e}}{M_{3}}.$ (121)
Hereafter, we redefine $\epsilon_{a}=\frac{a_{e}}{M_{1}}$, $\epsilon_{b}=\frac{b_{e}}{M_{2}}$, and $\epsilon_{c}=\frac{c_{e}}{M_{3}}$.
Then, the neutrino mass matrix is expanded via $\epsilon_{a}$ in the limit of $\epsilon=0$, and its form is given by
$m_{\nu}^{0}\simeq b_{0}^{2}\frac{v_{\Delta}y_{\Delta}}{9\sqrt{2}}\left[\tilde{y}^{\prime}_{\Delta}\epsilon_{b}\epsilon_{c}\left(\begin{array}[]{ccc}-4&2\omega^{2}&5\omega\\ 2\omega^{2}&8\omega&2\\ 5\omega&2&-4\omega^{2}\end{array}\right)+\left(\begin{array}[]{ccc}4&-2\omega^{2}&4\omega\\ -2\omega^{2}&\omega&-2\\ 4\omega&-2&4\omega^{2}\end{array}\right)\epsilon_{a}^{2}\right].$ (128)
The mixing matrix is given by
$U_{\nu}^{0}=\frac{1}{3}\left(\begin{array}[]{ccc}2\omega&-\omega^{2}&2\\ -\frac{3\omega}{\sqrt{2}}&0&\frac{3}{\sqrt{2}}\\ \frac{\omega^{2}}{\sqrt{2}}&2\sqrt{2}&\frac{\omega}{\sqrt{2}}\end{array}\right).$ (132)
Then, the mass eigenvalues are found as
$D_{\nu_{1}}^{0}=b_{0}^{2}\epsilon_{a}^{2},\ D_{\nu_{2}}^{0}=b_{0}^{2}\tilde{y}^{\prime}_{\Delta}\epsilon_{b}\epsilon_{c},,\ D_{\nu_{3}}^{0}=b_{0}^{2}\tilde{y}^{\prime}_{\Delta}\epsilon_{b}\epsilon_{c}.$ (133)
As a result, we have two large mixing angles of PMNS at the leading order. Below, we will numerically show there exist solutions at nearby $\tau=\omega$.
IV Numerical analysis and phenomenology
In this section, we perform chi square numerical analysis to fit the neutrino data and discuss several features of our model where we refer to NuFit 5.2 Esteban et al. (2020). The $\Delta\chi^{2}$ value is calculated as
$\Delta\chi^{2}=\sum_{i}\left(\frac{O_{i}^{\rm obs}-O_{i}^{\rm th}}{\delta O_{i}^{\rm exp}}\right)^{2},$ (134)
[…]
In addition to modulus $\tau\sim\omega$, we have 13 relevant free parameters for lepton mass matrices that are $\{v^{\prime},v_{\Delta},M_{1},M_{2},M_{3},a_{e},a_{b},a_{c},f_{1},f_{2},f_{3},y_{\Delta},y^{\prime}_{\Delta}\}$ where only $y^{\prime}_{\Delta}$ is complex. Three parameters, $\{a_{e},b_{e},c_{e}\}$, are fixed to fit charged lepton masses via Eq. (39). Six parameters, $\{y_{\Delta},v_{\Delta},v^{\prime},M_{1},f_{1},y_{\Delta}\}$, are factored out and they are only related to overall scale of mass matrices. Then neutrino mass squared differences and mixings are fitted by remaining rescaled free parameters that are $\{\tilde{f}_{2},\tilde{f}_{3},\tilde{M}_{2},\tilde{M}_{3},\tilde{y}_{\Delta}\}$ where $\tilde{f}_{2,3}\equiv f_{2,3}/f_{1}$ and $\tilde{M}_{2,3}\equiv M_{2,3}/M_{1}$. We randomly select all the rescaled input parameters within the ranges of $10^{-3}-10^{3}]$ except for $|\tilde{y}^{\prime}_{\Delta}|$ that is $10^{-5}-1]$ through our numerical analysis. We take mass scale as $m^{\prime}>1$ TeV, $M_{1,2,3}>100$ TeV to satisfy our assumption regarding hierarchy of mass parameters and CLFV constraints discussed in the previous section. In addition, we implicitly satisfy $m^{\prime}/M_{1}\lesssim 10^{-3}$ to evade non-unitarity bounds.